OpenStudy (anonymous):

(x+3)^2 - (3x-1)^2 = x^3 + 4? i've tried this three times and I still get the wrong answer!

OpenStudy (anonymous):

\[(x+3)^2 - (3x - 1)^2 = x^3 + 4\]

OpenStudy (anonymous):

just to clarify

jimthompson5910 (jim_thompson5910):

unless you know the cubic formula, your best bet is to use a graphing calculator to find the approximate solutions.

OpenStudy (anonymous):

we're not allowed to use a calculator. the answer is supposed to come out to be - 2/3

OpenStudy (anonymous):

but I don't know how to get there

jimthompson5910 (jim_thompson5910):

ok so there's a typo somewhere in the problem then, my guess is that x^3 is one of the typos

OpenStudy (anonymous):

wait, I made a mistake typing out the problem. it's actually \[(x+3)^3\] as the first one

jimthompson5910 (jim_thompson5910):

oh ok thx, one sec

OpenStudy (anonymous):

so \[(x+3)^3 - (3x - 1)^2 = x^3 + 4\]

jimthompson5910 (jim_thompson5910):

\[\large (x+3)^3- (3x-1)^2 = x^3 + 4\] \[\large x^3+9x^2+27x+27- (9x^2-6x+1) = x^3 + 4\] \[\large x^3+9x^2+27x+27- 9x^2+6x-1 = x^3 + 4\] \[\large x^3+9x^2+27x+27- 9x^2+6x-1 - x^3 - 4=0\] \[\large 33x+22=0\] \[\large 33x=-22\] \[\large x=-\frac{22}{33}\] \[\large x=-\frac{2}{3}\]

OpenStudy (anonymous):

thank you! i see exactly where I went wrong.